OpenStudy (anonymous):

Transition Matrix - Markov Chains Let \[P _{2}=\left[\begin{matrix}1-\alpha & \alpha \\ \beta & 1-\beta\end{matrix}\right]\] Calculate it's n-step transition matrix \[(P_2)^n\] I know I need to start by finding the eigenvalues: So we set \[ \det(P - \lambda I)=0\] and I get that \[\alpha+\beta-1=0\] but how do I interpret my eigenvalues/vectors from this?

OpenStudy (experimentx):

how do you get that value? and what are your eiven values?

OpenStudy (anonymous):

I think I've just figured out exactly what I've done wrong!

OpenStudy (experimentx):

okay .. that's great!!

OpenStudy (anonymous):

I didn't actually use the formula, I just calculated the determinant and set it to zero!

OpenStudy (anonymous):

Actually no I haven't worked it out, so if I take the determinant and set it to zero, I get: \[(1-\alpha-\lambda)(1-\beta-\lambda)-\alpha \beta=0\] to which I got: \[\lambda^2+(\alpha+\beta-2)\lambda-(\alpha+\beta-1)=0\] So if:\[\alpha + \beta=0 \quad \text{we obtain an eigenvalue of 1}\] \[\text{if} \quad \alpha+\beta>0 \quad\text{then we obtain an eigenvalue of} \quad 1-\alpha-\beta\] How do I then calculate my eigenvectors?

OpenStudy (experimentx):

find your eiven values first ... use that eigen value in \[ P_2 X = \lambda X \] you will get two set of equations by comparing the rows of X. solve it and can you have your eigen vectors.